# CADS: A Systematic Literature Review on the Challenges of Abstractive Dialogue Summarization

### v2

@@ Abstract
Abstractive dialogue summarization is the task of distilling conversations into informative and concise summaries. Although focused reviews have been conducted on this topic, there is a lack of comprehensive work that details the core challenges of dialogue summarization, unifies the differing understanding of the task, and aligns proposed techniques, datasets, and evaluation metrics with the challenges. This article summarizes the research on Transformer-based abstractive summarization for English dialogues by systematically reviewing 1262 unique research papers published between 2019 and 2024, relying on the Semantic Scholar and DBLP databases. We cover the main challenges present in dialog summarization (i.e., language, structure, comprehension, speaker, salience, and factuality) and link them to corresponding techniques such as graph-based approaches, additional training tasks, and planning strategies, which typically overly rely on BART-based encoder-decoder models. We find that while some challenges, like language, have seen considerable progress, mainly due to training methods, others, such as comprehension, factuality, and salience, remain difficult and hold significant research opportunities. We further investigate how these approaches are typically assessed, covering the datasets for the subdomains of dialogue (e.g., meeting, customer service, and medical), the established automatic metrics (e.g., ROUGE), and common human evaluation approaches for assigning scores and evaluating annotator agreement. We observe that only a few datasets with community interest span across all subdomains. Despite its limitations, the ROUGE metric is the most commonly used, while human evaluation, considered the gold standard, is frequently reported without sufficient detail on inner-annotator agreement and annotation guidelines. Additionally, we discuss the possible implications of the recently explored large language models and conclude that despite a potential shift in relevance and difficulty, our described challenge taxonomy remains relevant.
@@ 1 Introduction
Figure 1: Overview of the big six challenges in dialogue summarization, a short description of the challenges, and the estimation of their progress derived from related sub-challenges. Green means mostly mitigated, orange means good progress, and red stands for marked challenges still exist.
@@
Abstractive dialogue summarization, a task within Natural Language Processing (NLP) and text summarization, entails condensing key information from conversations into succinct and coherent summaries (?). This sub-field of text summarization is gaining prominence and is relevant for various real-world scenarios, including customer service (e.g., social media (?) and e-commerce (?)), healthcare (?), daily life (?), meetings (?), and open-domain conversations (e.g., online-chat (?)). The relevance of this task, which offers insights into discussed concerns or issues, is thereby driven by the increase in digital conversations in these scenarios (?). While the common, manual approach to dialogue summarization is considered to be prone to human errors and consumes significant time and effort, automatic dialogue summarization may effectively reduce this overhead on the user side (?).
@@
Diverging from traditional text summarization, which typically focuses on formal and linear content such as news articles (?) or scientific publications (?), dialogue summarization confronts unique challenges (?): conversational text is inherently dynamic, interactive, and non-linear (?), marked by verbosity, repetition, and informality. Salient information is often distributed across multiple speakers, mixed with casual and off-topic remarks (?). Additionally, the use of elliptical and fragmented sentences, characterized by, e.g., incomplete thoughts and context-specific abbreviations, further hinders the summarization process. These complexities make traditional text summarization approaches less effective, which are not trained to bridge the gap between edited text and dialogue and struggle to transfer learned patterns through few-shot learning (?).
@@
Despite advancements achieved by leveraging and adapting general-purpose pre-trained neural language models (?, ?), and the recent trend to explore Large Language Models (LLMs) for this task (?), there remains a notable gap in the field’s foundational understanding of the challenges when processing conversation transcripts. While research papers introduce techniques addressing similar challenges, their definitions and interpretations of these challenges can vary markedly. For instance, while some describe the informal and ungrammatical nature of spoken language as a language challenge (?, ?), others frame the challenge as different individual styles and unstructured expressions (?, ?). Both points of view are thereby valid but, on their own, focus just on a subpart of the language challenge and blur the understanding of it. Current surveys leave out the challenges in dialogue summarization and their variations in definitions, focusing instead on details on datasets, metrics, and techniques. As a consequence, there is no clear understanding of which aspects of dialogue summarization are well-addressed and where significant research potential lies. Our systematic review consolidates the existing knowledge researched in the field considering Transformer-based models, aiming to provide a comprehensive overview of the challenges and their progress. We further consider the center interest in LLMs and discuss their impact on the relevance of the dialogue summarization challenges.
@@
This article focuses on four key areas, namely challenges, dataset overviews, techniques, and evaluation methods. In challenges, we categorize known challenges for Transformer-based models found when dealing with dialogue transcripts into six broader challenge blocks, i.e., language, structure, comprehension, speaker, salience, and factuality. Our proposed taxonomy is displayed in Figure 1, including the six challenge blocks, a short description, and exemplary sub-challenges of the challenge blocks. In datasets overview, our review displays used datasets focused on dialogue summarization, organized by their subdomain, and further an overview of techniques to generate datasets to cope with the existing data scarcity artificially and how to optimize data usage. In techniques of dialogue summarization, we then match 97 techniques proposed since 2019 to the six challenge blocks. Furthermore, in evaluation methods, we give an overview of typical evaluation metrics used, spanning count-based metrics (e.g., ROUGE (?)), model-based (e.g., BERTScore (?)), QA-based (e.g., QuestEval (?)), and human evaluation metrics for performance and annotator agreement. This systematic review contributes to a better understanding of the dialogue summarization problem, unifies the inherent definitions, overviews established techniques, and demonstrates the scarcity of datasets and fitting evaluation metrics.
@@
Organization of the Literature Review. The remainder of this review is structured as follows. We introduce the current state of dialogue summarization in Section 1.1 and show our methodology in Section 2. In Section 3, we provide the general problem definition and our challenges taxonomy consisting of six challenge blocks inherent in conversation scripts, further linking proposed approaches to handle the challenges. Section 4 overviews prominent datasets while the display of evaluation approaches in Section 5 helps readers choose suitable indices to evaluate the effectiveness of a model. Finally, Section 6 discusses future research directions followed by conclusions in Section 7. All resources for our review are publicly available. [^1] https://anonymous.4open.science/r/LitRev-DialogueSum/
@@ 1.1 Existing Reviews in Dialogue Summarization
Comprehensive literature reviews on dialogue summarization remain scarce, with notable contributions by ? (?) and ? (?). Existing surveys mainly explore the subdomains, their datasets, techniques, and metrics within dialogue summarization. They provide a high-level overview and categorization of prominent subdomains like meeting, email, chat, medical, and customer service but often neglect niche areas such as e-commerce and debates (?). On the datasets, ? (?) provide a framework linking linguistic dialogue types (?) (e.g., persuasion, negotiation, inquiry) with established datasets, offering insights into dataset suitability. ? (?) analyze dialogue summarization techniques up to early 2022, categorizing techniques by conversational context and component modeling (e.g., speaker, addressee, and utterance modeling to solve the ’Who says what to whom’ paradigm), covering a subset of the conversation challenges indirectly. ? (?) extend this analysis with techniques from 2023, exploring approaches like hierarchical models for long inputs (?), feature injection for enhanced understanding, auxiliary tasks for broader training objectives, and data augmentation from related tasks. Regarding evaluations, works on dialogue summarization leverage automatic metrics from traditional document summarization, with the study by ? (?) discussing their individual strengths and limitations.
While existing surveys provide an understanding of dialogue summarization regarding subdomains, datasets, techniques, and metrics, the field misses a link between these categories through a common framework, such as the challenges inherent in processing dialogues. Our review addresses this gap by introducing a comprehensive taxonomy of dialogue challenges and categorizing recent techniques until 2024 based on these challenges (Section 3). This approach highlights well-explored areas and points out gaps in current research. We analyze datasets (Section 4) and evaluation metrics (Section 5), linking them to the identified challenges. Additionally, we discuss the implications of the shift from smaller encoder-decoder models such as BART (?) to LLMs and the implications for the relevance of the challenges with this new backbone model type (Section 6). Our review concentrates exclusively on dialogue summarization and does not encompass broader areas such as general text summarization (?) or dialogue generation (?).
@@ 2 Methodology
We use the PRISMA checklist (?) for organizing our literature review, as it is an established approach to set up a comprehensive systematic literature review while reducing the potential for incomplete data and biases in content selection and presentation (?). This ensures that our approach is structured, transparent, and reproducible.
Figure 2: An overview of the different stages of our methodology and their inherent sub-stages.
Our methodology comprises two stages as detailed in Figure 2.
@@
We evaluate 1262 unique papers from 2019 until 2024, offering an up-to-date perspective on the state of the art in dialogue summarization using Transformer-based architectures. We decide against including references from selected papers that did not appear in our crawl as additional sources to maintain an unbiased selection process and avoid complicating the stopping criterion for literature inclusion. Considering our exclusion criteria, such as multilingualism, non-abstractive, and non-Transformer-based approaches, we select 133 papers for our literature review.
@@ 2.1 Stage 1: Retrieving Candidate Documents
Information Sources. Our primary data sources are Semantic Scholar [^2] https://www.semanticscholar.org/ and DBLP [^3] https://dblp.org/ (DataBase systems and Logic Programming), both recognized for extensive coverage in computer science (?, ?). While Semantic Scholar allows for advanced search functionalities and complex queries (?, ?, ?), DBLP is the most comprehensive database for computer science publications, frequently used in other surveys as sole source (?, ?, ?) and encompassing major peer-reviewed journals and conference proceedings. This dual-source strategy ensures thorough retrieval and coverage specific to dialogue summarization, strengthening the certainty and reliability of our evidence base. To ensure our literature review remains up-to-date, we regularly re-crawl databases while conducting the review to incorporate the latest publications into our analysis.  [^4] Last crawled on March 25, 2024. While we acknowledge the potential for these databases to miss the most recent uploads, our strategy minimizes this risk. We exclude databases such as Google Scholar, which overly returned grey literature, and Web of Science, which emphasized extractive summarization techniques during our pre-testing, to maintain a focus on peer-reviewed literature in abstractive dialogue summarization. We design an automatic pipeline to retrieve query-related works and remove duplicates. We limit the number of results per query to the top 100 based on the relevance ranking provided by the platforms, which was rarely required. We recognize that the ranking systems may be biased towards prestigious journals or highly cited authors. Hence, we retrieve all papers published after June 2023, not relying on the ranking systems to account for their limited time to gather citations.
Data Collection Process. Our search strategy uses key terms related to dialogue summarization identified during our initial related work analysis (Section 1.1). Anchor terms such as ’dialogue summarization’ and ’multi-party conversation’ are combined with specific search terms such as ’technique’. We use four anchor terms and 28 search terms, resulting in 112 search queries for each database. We present the complete list and search keywords in our repository [^5] https://anonymous.4open.science/r/LitRev-DialogueSum/ . Employing a Python-based pipeline (Section 8), we systematically extract documents from Semantic Scholar and DBLP APIs, merging and unifying the results into tabular data. Between January 2019 and March 2024, we retrieved 732 publications from DBLP and 968 from Semantic Scholar. After removing 438 duplicates identified across both sources, we have 1262 unique publications at the end of the stage. These results are tagged with their respective queries and compiled into a CSV file for subsequent selection and analysis.
@@ 2.2 Stage 2: Manual Filtering
Eligibility Criteria. For our manual filtering process, we define eligibility criteria to ensure the focus of our literature review. We look at long and short, open-access papers published since 2019 that align with the rise of Transformer-based text summarization models such as BART (?) and PEGASUS (?), which have been the core model for most approaches in the field. This timeframe further allows us to concentrate on current challenges rather than including some that may have lost relevance (e.g., negation (?)) because of new advancements in the field. We exclude non-English primary datasets, multi-modal studies, and extractive or non-Transformer-based methods to maintain a focus on English-language research on abstractive dialogue summarization.
Manual Paper Processing. A team of three reviewers  [^6] Each team member discloses potential conflicts of interest to ensure an unbiased review process., consisting of postgraduate students and doctoral candidates with a computer science background, undertake the manual screening process, starting with titles and abstracts to narrow down the automatic selection from 1262 to 186 documents using our eligibility criteria. They then review the full texts, checking if the works still meet the eligibility criteria, and decide whether to include, exclude, or discuss each paper. Disagreements are resolved through consensus. They also track features such as used datasets, employed backbone models, proposed techniques, and evaluation approaches. This step filters out 53 papers, leaving 133 to compose the literature review and detail challenges, techniques, datasets, and evaluation metrics. To further minimize bias, a secondary review is performed as needed, reaffirming initial selections and categorizations, thereby enhancing the study’s integrity.
@@ 3 Challenges and Progress
Abstractive dialogue summarization aims to condense multi-party conversations into their key points, ranging from casual chats to expert discussions, and whether these points were explicitly stated or inferred (?). The task was first thoroughly defined in 2002 by ? (?), outlining challenges unique to spoken dialogues such as handling speech disfluency, the lack of clear sentence boundaries, distribution of salient information across various speakers and turns, resolving references, identifying discourse structures, and dealing with inaccuracies caused by speech recognition errors. While some definitions take a mathematical approach (?), viewing dialogues as sequences of turns that are compressed and each associated with a specific speaker, others adopt a broader description of the task as ”extracting useful information from a dialogue” (?). Although all definitions contain relevant, overlapping aspects, they lay down different viewpoints, each considering a different set of challenges to define dialogue summarization. Our goal is to organize these definitions and provide a concise, structured overview of the progress in abstractive dialogue summarization, extending the earlier challenge-based concept (?) and proposing the Challenges of Abstractive Dialogue Summarization (CADS) Taxonomy (Figure 1) tailored to the current Transformer-based model architecture. This taxonomy consists of six challenges: language, structure, comprehension, speaker, salience, and factuality. The first five challenges are thereby related to the input, while the last one concerns reliability. In the following Sections 3.1, 3.2, 3.3, 3.4, 3.5 and 3.6, we detail the CADS Taxonomy as shown in Figure 1.
For this section, we consider 91 papers in total, from which 73 papers define the challenge characteristics (sub-challenges in italic) and 74 describe explored techniques. We exclude papers mainly discussing evaluation strategies and datasets from this subset and cover them in Sections 4 and 5.

### v3

@@ Abstract
Abstractive dialogue summarization is the task of distilling conversations into informative and concise summaries. Although focused reviews have been conducted on this topic, there is a lack of comprehensive work that details the core challenges of dialogue summarization, unifies the differing understanding of the task, and aligns proposed techniques, datasets, and evaluation metrics with the challenges. This article summarizes the research on Transformer-based abstractive summarization for English dialogues by systematically reviewing 1262 unique research papers published between 2019 and 2024, relying on the Semantic Scholar and DBLP databases. We cover the main challenges present in dialog summarization (i.e., language, structure, comprehension, speaker, salience, and factuality) and link them to corresponding techniques such as graph-based approaches, additional training tasks, and planning strategies, which typically overly rely on BART-based encoder-decoder models. Recent advances in training methods have led to substantial improvements in language-related challenges. However, challenges such as comprehension, factuality, and salience remain difficult and present significant research opportunities. We further investigate how these approaches are typically analyzed, covering the datasets for the subdomains of dialogue (e.g., meeting, customer service, and medical), the established automatic metrics (e.g., ROUGE), and common human evaluation approaches for assigning scores and evaluating annotator agreement. We observe that only a few datasets (i.e., SAMSum, AMI, DialogSum) are widely used. Despite its limitations, the ROUGE metric is the most commonly used, while human evaluation, considered the gold standard, is frequently reported without sufficient detail on the inter-annotator agreement and annotation guidelines. Additionally, we discuss the possible implications of the recently explored large language models and conclude that our described challenge taxonomy remains relevant despite a potential shift in relevance and difficulty.
@@ 1 Introduction
Figure 1: Overview of the six challenges in dialogue summarization, including a brief description of each challenge and an estimation of progress for related sub-challenges. Progress is evaluated based on two factors: (1) the extent to which mitigation strategies have been developed, and (2) the measurable improvement in summary quality as a result of these strategies. Green means mostly mitigated, orange means good progress, and red stands for marked challenges still exist.
@@
Abstractive dialogue summarization, a task within Natural Language Processing (NLP) and text summarization, entails condensing key information from conversations into succinct and coherent summaries (?). This sub-field of text summarization is gaining prominence and is relevant for various real-world scenarios, including customer service (e.g., social media, ?, 2021, and e-commerce, ?, 2022), healthcare (?), daily life (?), meetings (?), and open-domain conversations (e.g., online-chat, ?, 2019). The relevance of this task, which offers insights into discussed concerns or issues, is thereby driven by the increase in digital conversations in these scenarios (?). While the common, manual approach to dialogue summarization is considered to be prone to human errors and consumes significant time and effort, automatic dialogue summarization may effectively reduce this overhead on the user side (?).
@@
Diverging from traditional text summarization, which typically focuses on formal and linear content such as news articles (?) or scientific publications (?), dialogue summarization demonstrates unique challenges (?): conversational text is inherently dynamic, interactive, and non-linear (?), marked by verbosity, repetition, and informality. Salient information is often distributed across multiple speakers, mixed with casual and off-topic remarks (?). Additionally, the use of elliptical and fragmented sentences, characterized by, e.g., incomplete thoughts and context-specific abbreviations, further hinders the summarization process. These complexities make traditional text summarization approaches less effective, as they are not trained to bridge the gap between edited text and dialogue and struggle to transfer learned patterns through few-shot learning (?).
@@
Despite advancements achieved by leveraging and adapting general-purpose pre-trained neural language models (?, ?), and the recent trend to explore Large Language Models (LLMs) for this task (?), there remains a notable gap in the field’s foundational understanding of the challenges when processing conversation transcripts. While research papers introduce techniques addressing similar challenges, their definitions and interpretations of these challenges can vary markedly. For instance, while some describe the informal and ungrammatical nature of spoken language as a language challenge (?, ?), others frame the language challenge as different individual styles and unstructured expressions (?, ?). Both viewpoints are thereby valid but, on their own, focus just on a subpart of the language challenge and blur the complete understanding of it. Current surveys leave out the challenges in dialogue summarization and their variations in definitions, focusing instead on details on datasets, metrics, and techniques. Consequently, there is no clear understanding of which aspects of dialogue summarization are well-addressed and where significant research potential lies. Our systematic review consolidates the knowledge researched in the field considering Transformer-based models, aiming to provide a comprehensive overview of the challenges and their progress. We further consider the interest in LLMs and discuss their impact on the relevance of the dialogue summarization challenges.
@@
This article focuses on four key areas, namely challenges, dataset overviews, techniques, and evaluation methods. In challenges, we categorize known hurdles for Transformer-based models found when dealing with dialogue transcripts into six broader challenge blocks, i.e., language, structure, comprehension, speaker, salience, and factuality. Our proposed taxonomy is displayed in Figure 1, including the six challenge blocks, a short description, and exemplary sub-challenges of the challenge blocks. In datasets overview, our review displays used datasets focused on dialogue summarization, organized by their subdomain, and further an overview of techniques to generate datasets to cope with the existing data scarcity artificially and how to optimize data usage. In techniques of dialogue summarization, we then match 97 techniques proposed since 2019 to the six challenge blocks. Furthermore, in evaluation methods, we provide an overview of typical evaluation metrics used, spanning count-based metrics (e.g., ROUGE, ?, 2004), model-based (e.g., BERTScore, ?, 2020), QA-based (e.g., QuestEval, ?, 2021), and human evaluation metrics for performance and annotator agreement. This systematic review contributes to a better understanding of the dialogue summarization problem, unifies the inherent definitions, overviews established techniques, and demonstrates the scarcity of datasets and fitting evaluation metrics.
@@
Organization of the Literature Review. The remainder of this review is structured as follows. We introduce the current state of dialogue summarization in Section 1.1 and show our methodology in Section 2. In Section 3, we provide the general problem definition and our challenges taxonomy consisting of six challenge blocks inherent in conversation scripts, further linking proposed approaches to handle the challenges. Section 4 overviews prominent datasets, while the display of evaluation approaches in Section 5 helps readers choose suitable indices to evaluate the effectiveness of a model. Finally, Section 6 discusses future research directions followed by conclusions in Section 7. All resources for our review are publicly available. [^1] https://github.com/FKIRSTE/LitRev-DialogueSum
@@ 1.1 Existing Reviews in Dialogue Summarization
Comprehensive literature reviews on dialogue summarization remain scarce, with notable contributions by ? (?) and ? (?). Existing surveys mainly explore the subdomains, their datasets, techniques, and metrics within dialogue summarization. They provide a high-level overview and categorization of prominent subdomains like meeting, email, chat, medical, and customer service but often neglect niche areas such as e-commerce and debates (?). On the datasets, ? (?) provide a framework linking linguistic dialogue types (?) (e.g., persuasion, negotiation, inquiry) with established datasets, offering insights into dataset suitability. ? (?) analyze dialogue summarization techniques up to early 2022, categorizing techniques by conversational context and component modeling (e.g., speaker, addressee, and utterance modeling to solve the ‘Who says what to whom’ paradigm), covering a subset of the conversation challenges indirectly. ? (?) extend this analysis with techniques from 2023, exploring approaches like hierarchical models for long inputs (?), feature injection for enhanced understanding, auxiliary tasks for broader training objectives, and data augmentation from related tasks. Regarding evaluations, works on dialogue summarization leverage automatic metrics from traditional document summarization, with the study by ? (?) discussing their individual strengths and limitations.
While existing surveys provide an understanding of dialogue summarization regarding subdomains, datasets, techniques, and metrics, the field misses a link between these categories through a common framework, such as the challenges inherent in processing dialogues. Our review addresses this gap by introducing a comprehensive taxonomy of dialogue challenges and categorizing recent techniques until 2024 based on these challenges (Section 3). This approach highlights well-explored areas and points out gaps in current research. We analyze datasets (Section 4) and evaluation metrics (Section 5), linking them to the identified challenges. Additionally, we discuss the shift from smaller encoder-decoder models such as BART (?) to LLMs and the implications for the relevance of the challenges with this new backbone model type (Section 6). Our review focuses exclusively on dialogue summarization and does not encompass broader areas such as general text summarization (?) or dialogue generation (?).
@@ 2 Methodology
We use the PRISMA checklist (?) for organizing our literature review, as it is an established approach to set up a comprehensive systematic literature review while reducing the potential for incomplete data and biases in content selection and presentation (?). This checklist ensures that our approach is structured, transparent, and reproducible.
Figure 2: An overview of the different stages of our methodology and their inherent sub-stages. The red boxes indicate the number of papers considered in the respective step.
Our methodology comprises two stages, as detailed in Figure 2.
@@
We evaluate 1262 unique papers from 2019 until 2024, offering an up-to-date perspective on the state of the art in dialogue summarization using Transformer-based architectures. We decided against including references from selected papers that did not appear in our crawl as additional sources to maintain an unbiased selection process and avoid complicating the stopping criterion for literature inclusion. Considering our exclusion criteria, such as multilingualism, non-abstractive, and non-Transformer-based approaches, we select 133 papers for our literature review.
@@ 2.1 Stage 1: Retrieving Candidate Documents
Information Sources. Our primary data sources are Semantic Scholar [^2] https://www.semanticscholar.org/ and DBLP [^3] https://dblp.org/ (DataBase systems and Logic Programming), both recognized for extensive coverage in computer science (?, ?). While Semantic Scholar allows for advanced search functionalities and complex queries (?, ?, ?), DBLP is the most comprehensive database for computer science publications, frequently used in other surveys as sole source (?, ?, ?) and encompassing major peer-reviewed journals and conference proceedings. This dual-source strategy ensures thorough retrieval and coverage specific to dialogue summarization, strengthening the certainty and reliability of our evidence base. To ensure our literature review remains up-to-date, we regularly re-crawl databases while conducting the review to incorporate the latest publications into our analysis. [^4] Last crawled on March 25, 2024. While we acknowledge the potential for these databases to miss the most recent uploads, our strategy minimizes this risk. We exclude databases such as Google Scholar, which overly returned gray literature, and Web of Science, which emphasized extractive summarization techniques during our pre-testing, to maintain a focus on peer-reviewed literature in abstractive dialogue summarization. We design an automatic pipeline to retrieve query-related works and remove duplicates. We limit the number of results per query to the top 100 based on the relevance ranking provided by the platforms, which was rarely required. We recognize that the ranking systems may be biased towards prestigious journals or highly cited authors. Hence, we retrieve all papers published after June 2023, not relying on the ranking systems to account for their limited time to gather citations.
Data Collection Process. Our search strategy uses key terms related to dialogue summarization identified during our initial related work analysis (Section 1.1). Anchor terms such as ‘dialogue summarization’ and ‘multi-party conversation’ are combined with specific search terms such as ‘technique.’ We use four anchor terms and 28 search terms, resulting in 112 search queries for each database. We present the complete list and search keywords in our repository [^5] https://github.com/FKIRSTE/LitRev-DialogueSum . Employing a Python-based pipeline (Section 9), we systematically extract documents from Semantic Scholar and DBLP APIs, merging and unifying the results into tabular data. Between January 2019 and March 2024, we retrieved 732 publications from DBLP and 968 from Semantic Scholar. After removing 438 duplicates identified across both sources, we conclude the stage with 1262 unique publications. These results are tagged with their respective queries and compiled into a CSV file for subsequent selection and analysis.
@@ 2.2 Stage 2: Manual Filtering
Eligibility Criteria. For our manual filtering process, we define eligibility criteria to ensure the focus of our literature review. We consider open-access papers published since 2019 that align with the rise of Transformer-based text summarization models such as BART (?) and PEGASUS (?), which have been the core model for most approaches in the field. This timeframe further allows us to concentrate on current challenges rather than including some that may have lost relevance (e.g., negation, ?, 2021) because of advancements in the field. We exclude non-English primary datasets, multi-modal studies, and extractive or non-Transformer-based methods and focus on English-language research on abstractive dialogue summarization.
Manual Paper Processing. A team of three reviewers [^6] Each team member discloses potential conflicts of interest to ensure an unbiased review process., consisting of postgraduate students and doctoral candidates with a computer science background, undertake the manual screening process, starting with titles and abstracts to narrow down the automatic selection from 1262 to 186 documents using our eligibility criteria. They then review the full texts, checking if the works still meet the eligibility criteria, and decide whether to include, exclude, or discuss each paper. Each paper undergoes a secondary review to minimize bias, reaffirming initial selections and categorizations, thereby enhancing the study’s integrity. Eventual disagreements are resolved through consensus. The annotators also track features such as used datasets, employed backbone models, proposed techniques, and evaluation approaches. This step filters out 53 papers, leaving 133 to compose the literature review and detail challenges, techniques, datasets, and evaluation metrics.
@@ 3 Challenges and Progress
Abstractive dialogue summarization aims to condense multi-party conversations into their key points, ranging from casual chats to expert discussions, and whether these points were explicitly stated or inferred (?). The task was first thoroughly defined in 2002 by ?, outlining challenges unique to spoken dialogues such as handling speech disfluency, the lack of clear sentence boundaries, distribution of salient information across various speakers and turns, resolving references, identifying discourse structures, and dealing with inaccuracies caused by speech recognition errors. While some definitions take a mathematical approach (?), viewing dialogues as sequences of turns that are compressed and associated with a specific speaker, other definitions adopt a broader description of the task as ‘extracting useful information from a dialogue’ (?). Although all definitions contain relevant, overlapping aspects, they have different viewpoints, each considering a different set of challenges to define dialogue summarization. Our goal is to organize these definitions and provide a concise, structured overview of the progress in abstractive dialogue summarization, extending the earlier challenge-based concept (?) and proposing the Challenges of Abstractive Dialogue Summarization (CADS) Taxonomy (Figure 1) tailored to the current Transformer-based model architecture. This taxonomy comprises six challenges: language, structure, comprehension, speaker, salience, and factuality. The first five challenges are related to the input, while the last concerns reliability. To develop this comprehensive framework, we employed human annotators (Section 2.2) to systematically analyze and categorize recurring challenges in the literature and consolidate them into the six challenges. Each pillar encompasses related sub-challenges that represent key focus areas consistently addressed in the literature along the broader challenges. In the following Sections 3.1, 3.2, 3.3, 3.4, 3.5 and 3.6, we detail the CADS Taxonomy as shown in Figure 1. An overview of the approach strategies proposed to mitigate the individual challenges is shown in Table 1. Table 2 summarizes the expected errors when challenges are not successfully mitigated, using the analysis by ? (?) as a base for the individual error types.
For this section, we consider 91 papers, from which 73 papers define the challenge characteristics (sub-challenges in italic) and 74 describe explored techniques. We exclude papers mainly discussing evaluation strategies and datasets from this subset and cover them in Sections 4 and 5.
Challenge Approach Category Related Papers
Language Pre-training ? (?, ?, ?, ?)
Training tasks ? (?, ?, ?, ?, ?)
Pre-processing ? (?)
Structure Pre-training ? (?, ?, ?)
Training tasks ? (?, ?, ?, ?)
Architecture modification ? (?, ?, ?, ?, ?)
Importance measures ? (?, ?)
Comprehension Architecture modification ? (?)
Speaker Training tasks ? (?, ?, ?, ?)
Architecture modification ? (?, ?, ?, ?)
Pre-processing ? (?, ?)
Post-processing ? (?, ?)
Salience Pre-training ? (?, ?)
Training tasks ? (?, ?, ?)
Architecture modification ? (?, ?)
Human feedback ? (?)
Loss function ? (?)
Pre-processing ? (?, ?)
Factuality Training tasks ? (?, ?)
Architecture modification ? (?, ?, ?, ?)
Human feedback ? (?)
Loss function ? (?, ?)
Post-processing ? (?, ?)
Table 1: Overview of challenges, major approach trends, and corresponding literature. More details are stated in the article’s accompanying repository:
https://github.com/FKIRSTE/LitRev-DialogueSum.